five years ago, Julia was twice as old as Karla. five years from now , karla's age will be three-forths of julia's age. How old is each girl now?
Answer by ptaylor(1894)

Let x=Julia's age now
x-5=Julia's age five years ago
x+5=Julia's age five years from now
And let y=Karla's age now
y-5=Karla's age five years ago
y+5=Karla's five years from now
Now we are told the following:
x-5=2(y-5)---------------------------eq1
y+5=(3/4)(x+5)----------------------------eq2
simplify eq1:
x-5=2y-10 add 5 to both sides
x=2y-5----------------------eq1
simplify eq2:
4y+20=3x+15 subtract 20 from both sides
4y=3x-5------------------------eq2
substitute x=2y-5 into eq2
4y=3(2y-5)-5 get rid of parens
4y=6y-15-5 subtract 6y from both sides
-2y=-20 divide both sides by -2
y=10----------------------Karla's age now
x=2y-5
x=2*10-5
x=15--------------------------Julia's age now
CK
Five years ago:
Karla was 5 and Julia was 10
Julia was twice as old as Karla
Five years from now:
Karla will be 15 and Julia will be 20
15 is 3/4 of 20
Hope this helps----ptaylor